# sample.bib Deep Learning and LLM-based Methods Applied to Stellar Lightcurve Classification

### v1

@@ Abstract
Light curves serve as a valuable source of information on stellar formation and evolution. With the rapid advancement of machine learning techniques, it can be effectively processed to extract astronomical patterns and information. In this study, we present a comprehensive evaluation of deep-learning and large language model (LLM) based models for the automatic classification of variable star light curves, based on large datasets from the Kepler and K2 missions. Special emphasis is placed on Cepheids, RR Lyrae, and eclipsing binaries, examining the influence of observational cadence and phase distribution on classification precision. Employing AutoDL optimization, we achieve striking performance with the 1D-Convolution+BiLSTM architecture and the Swin Transformer, hitting accuracies of 94% and 99% correspondingly, with the latter demonstrating a notable 83% accuracy in discerning the elusive Type II Cepheids—comprising merely 0.02% of the total dataset. We unveil StarWhisper LightCurve (LC), an innovative Series comprising three LLM-based models: LLM, multimodal large language model (MLLM), and Large Audio Language Model (LALM). Each model is fine-tuned with strategic prompt engineering and customized training methods to explore the emergent abilities of these models for astronomical data. Remarkably, StarWhisper LC Series exhibit high accuracies around 90%, significantly reducing the need for explicit feature engineering, thereby paving the way for streamlined parallel data processing and the progression of multifaceted multimodal models in astronomical applications. The study furnishes two detailed catalogs illustrating the impacts of phase and sampling intervals on deep learning classification accuracy, showing that a substantial decrease of up to 14% in observation duration and 21% in sampling points can be realized without compromising accuracy by more than 10%.
1 Introduction
@@
The phenomenon of light variation is a crucial aspect of astrophysics and has long been studied in the field of time-domain astronomy. Cepheid variables, a special kind of variable star, serve a critical role as a universal standard candle, enabling us to measure the distance to clusters and galaxies with their period-luminosity relation. Other types of variable stars, including RR Lyrae, Sct, and Dor stars, are characterized by their unique absolute magnitudes, pulsation patterns, each exhibiting diverse period-luminosity relations. The classification of these pulsating stars has significantly enriched our understanding of their formation and evolution. Moreover, it has shed light on the structure and dynamics of binary, n-body systems and our Galaxy. [prsaKEPLERECLIPSINGBINARY2011, slawsonKEPLERECLIPSINGBINARY2011, matijevicKEPLERECLIPSINGBINARY2012, conroyKEPLERECLIPSINGBINARY2014a, conroyKeplerEclipsingBinary2014b, lacourseKeplerEclipsingBinary2015, kirkKEPLERECLIPSINGBINARY2016, abdul-masihKEPLERECLIPSINGBINARY2016a]
@@
Long periods of observations are crucial in order to understand the nature of variable stars. To carry out these observations, it is important to make wise use of telescope time that is a valuable resource [garcia-piquerEfficientSchedulingAstronomical2017]. One such strategy involves the use of simulations to create plans that optimally allocate available telescope time, considering overheads and other factors. Another strategy employs autonomous agents to optimize the observation of time-varying phenomena [saundersOptimalObservingAstronomical]. In this study, we aim to understand the importance tied to different phases of variables and the classification accuracy provided by varying cadences. This can help us to establish robust guidelines for future star gazing observations [wangTransferLearningApplied2023].
@@
A substantial amount of labeled data is required, in order to understand the significance of phase and cadence for different variable stars. In recent years, astronomy has entered the era of big data. For example, the Zwicky Transient Factory (ZTF) [bellmZwickyTransientFacility2014] at Palomar Observatory scans the sky every two days with a 1.2m telescope , generating 1TB of raw data each night [mahabalMachineLearningZwicky2019a]. The upcoming SiTian survey [liuSiTianProject2021], aims to monitor over 1,000 of sky every 30 minutes using fifty 1m class Schmidt telescopes, and is expected to produce around 140TB of processed data each night.
@@
As the volume of data grows, there’s a increasing requirement for efficient and automated interpretation and analysis methods. Deep learning, a component of machine learning, has risen as a powerful tool for image and signal processing [lecunDeepLearning2015]. It can learn and extract features from data [krizhevskyImageNetClassificationDeep2012a, lecunDeepLearning2015], making it an ideal solution for managing large, complex datasets. Specifically, recurrent neural networks (RNNs) have shown great effectiveness in processing time series data [liptonCriticalReviewRecurrent2015], while convolutional neural networks (CNNs) have demonstrated their superiority in image processing tasks [krizhevskyImageNetClassificationDeep2012a, heDeepResidualLearning2015]. Moreover, the transformer architecture, with its attention mechanism [vaswaniAttentionAllYou2017], has shown remarkable performance across various applications [devlinBERTPretrainingDeep2019a, radfordImprovingLanguageUnderstanding, radfordLanguageModelsAre, brownLanguageModelsAre2020]. RNNs originally developed by [elmanFindingStructureTime1990], led to Simple RNNs, which were later enhanced by LSTMs [hochreiterLongShortTermMemory1997] to address long-term dependencies. These were further simplified by GRUs [choPropertiesNeuralMachine2014] for computational efficiency. CNNs initially introduced for hand-written digit recognition [lecunBackpropagationAppliedHandwritten1989], and have evolved with architectures like AlexNet, VGGNet, and ResNet. Recent advancements include EfficientNet [tanEfficientNetRethinkingModel2020, tanEfficientNetV2SmallerModels2021], which optimizes depth, width, and resolution for model accuracy and efficiency.
@@
In recent years, Transformer architecture has shown remarkable performance across a diverse range of applications.Initially introduced to address sequential data processing tasks [vaswaniAttentionAllYou2017]. Transformer use self-attention mechanisms to model relationships with sequences, enabling them to capture long-range dependencies and relationships between sequence elements. The success of the Transformer in natural language processing has sparked interest in their potential application in other domains, such as computer vision and speech recognition. An example of a vision Transformer is the Swin Transformer [liuSwinTransformerHierarchical2021, liuSwinTransformerV22022], which has optimized the computation of attention mechanism and shown promising results in various benchmark datasets.
@@
Beyond the traditional reliance on extensive data for transfer learning, it is imperative to investigate the potential of LLM that incorporate both substantial datasets and large parameter counts. StarWhisper 111https://github.com/Yu-Yang-Li/StarWhisper is a LLM for astronomy, which has strong astronomical ability and instruction following ability, and can complete a series of functions such as knowledge question answering, calling multi-modal tools, and docking telescope control systems. The StarWhisper LC Series is initiated with the purpose of leveraging the experience gained from the prior training of the StarWhisper language model. It seeks to explore and discuss the potential harnessed from vast data to engender emergent properties in the analysis of light curve data. LLMs, such as the Gemini 7B model, have shown promise in adapting to new data types through fine-tuning with specific prompt templates [zhou2023fits13]. MLLMs, like the deepseek-vl-7b-chat, are adept at handling tasks involving image classification due to their extensive training on datasets containing chart data [tsai2019multimodal_6]. LALMs, such as Qwen-audio,trained on audio datasets, exhibits exceptional performance in audio classification. [yang2022voice2series_4].
@@
In this study, we perform a comprehensive evaluation of deep-learning and LLM-based models for the classification of variable star light curves, using the data from Kepler and K2 observations. The types of variable stars and the data pre-processing are presented in Section 2. In Section 3, we utilize various classification models, including the LSTM, the GRU, the Transformer, the LightGBM, the EfficientNet, the Swin Transformer and StarWhisper LC serires. Training methods are introduced in Section 4. We also present model performance and catalogs of the phase importance and sampling intervals in Section 5. A discussion is provided in Section 6.
2 Data
2.1 Kepler and K2
@@
The Kepler spacecraft was launched in 2009 aiming to discover for Earth-like planets [doi:10.1126/science.1185402]. It was equipped with an optical telescope a 95cm aperture and a field of view. The advanced technology allowed Kepler to precisely track light curves from 200,000 different targets. This high precision resulted in the discovery of over 2,000 planets. For stars with V band magnitude between 13mag to 14mag, the precision was 100 ppm (parts per million), while for stars with V band magnitude between 9mag to 10mag, the precision was 10 ppm. Unfortunately, after four years, half of Kepler’s four reaction wheels failed, leading to the end of its primary mission. Yet, this marked the beginning of the K2 mission. The K2 mission used the transit method to detect changes in light along the ecliptic plane and created catalogs with photometric precision closely matching that of the original Kepler mission. The observations of K2 were controlled using the remaining reaction wheels and thrusters, with each campaign limited to 80 days.
@@
There are two types of light curve available, named Presearch Data Conditioning (PDC) and Simple Aperture Photometry (SAP) light curves [Cleve2016KeplerIH]. While SAP light curves retain long-term trends, PDC light curves were generated by the Kepler Operation Science Center, and are free from systematic errors. Therefore, we adopt PDC light curves for our analysis. Additionally, we consider two types of data with different time resolutions: long-cadence data, with a 30-minute sampling interval, and short-cadence data, with a one-minute sampling interval. Due to the limited number of short-cadence data, we focus on the long-cadence light curves. For each light curve, we remove the quarter-to-quarter differences and convert the flux into relative flux, similar to the methods used in [yangFlareCatalogFlare2019, hanStellarActivityCycles2021].
2.2 Variable Stars
@@
Our training samples are similar to [wangTransferLearningApplied2023], including eclipsing binaries, RR Lyrae, Scuti, Dor, and Scuti / Dor hybrids. The type II Cepheids are included, in order to make a more universal sample. Table 1 lists our training sample and corresponding references. The training samples are seriously biased among different variable types. This bias often leads to decreased performance, especially in situations few-shot or small sample scenarios [kokolMachineLearningSmall2022, clemenconStatisticalLearningBiased2019]. However, [taniguchiMachineLearningModel2018] suggest that certain models may help overcome these challenges, although such cases seem to be the exception rather than the rule. These biased training samples offer a unique opportunity to study the algorithm’s dependence on such biases and to understand how this dependence affects the overall performance and accuracy when applied to astronomical data.
2.3 Pre-processing
@@
A pre-processing method, as described by [wangTransferLearningApplied2023], was adopted to manage and clean the light curves, with the aim of enhancing their features and expanding the training data. The lightcurves were segmented into 10-day intervals. Any segments with gaps exceeding a day were removed, while those with gaps less than one day were interpolated with a time sequence of 30 minute intervals.
| Label | Input Catalog | Final Sample | References |
| Sct | 1389 | 111528 | (1), (2), (3), (4), (14), (15) |
| EB | 2908 | 226937 | (6), (7), (8), (9), (10), (11), (12), (13) |
| Dor | 941 | 65786 | (14),(15) |
| HYB | 1552 | 33751 | (1), (2), (3), (4), (14), (15) |
| RR | 482 | 9306 | (5) |
| T2CEP | 3 | 94 | (1), (2), (3), (4), (5) |
| Total | 7275 | 447402 | (1)-(15) |
@@
The input catalog column indicates the sources of data that we obtained from several catalogs. The final sample column represents the final sample size after applying all the pre-processing procedures. References: (1)[kholopovCombinedGeneralCatalogue1998], (2)[durlevichListErrorsGCVS1994], (3)[artyukhinaVizieROnlineData1996], (4)[samus84thNameListVariable2021], (5)[molnarGaiaDataRelease2018], (6)[prsaKEPLERECLIPSINGBINARY2011], (7)[slawsonKEPLERECLIPSINGBINARY2011], (8)[matijevicKEPLERECLIPSINGBINARY2012], (9)[conroyKEPLERECLIPSINGBINARY2014a], (10)[conroyKeplerEclipsingBinary2014b], (11)[lacourseKeplerEclipsingBinary2015], (12)[kirkKEPLERECLIPSINGBINARY2016], (13)[abdul-masihKEPLERECLIPSINGBINARY2016], (14)[bradleyRESULTSSEARCHDOR2015], (15)[balonaGaiaLuminositiesPulsating2018].
@@
The table shows the period of the stars (in the unit of days) along with the variation of accuracy and time saved (percentages), for both phase importance and sampling research.
An alternative approach to time-series data classification involves categorizing of images that are created from graphically represented light curves. This is achieved using transfer learning techniques. As highlighted in their result, the continuous wavelet transform (CWT) method has demonstrated superior results in imaging light curves.
The Morlet wavelet was chosen as the core function for the CWT, owing to its proven effectiveness in analyzing signals that display shifts in amplitude. This analysis resulted in a collection of images that provide valuable insights into the time-frequency characteristics of the signals. Importantly, these images facilitate the identification of patterns, trends, and anomalies, thus offering a comprehensive understanding of the signals. Figure 1 shows some variable stars from the training sample.
2.4 Lomb-Scargle Periodogram
We employ the Lomb-Scargle algorithm to extract the most significant periods complete light curves, which are used to study the phase importance of periodic variables. The Lomb-Scargle algorithm, a variant of the Discrete Fourier Transform (DFT) developed by [1976Ap&SS..39..447L] and [scargleStudiesAstronomicalTime1982], has been specifically designed for unevenly sampled time-series data. It transforms a time series into a linear combination of sinusoidal waveforms, simplifying the conversion from time to frequency domain. We calculate the period by applying the Lightkurve Collaboration’s LSP method [vanderplasUnderstandingLombScarglePeriodogram2018]. Some examples are shown in Figure 2.
@@ 2.4 Lomb-Scargle Periodogram
3 Model Construction
In order to classify variable stars from their light curves, we have explored various deep learning models. The basic architectures of these models are shown in Figure 3.
3.1 CNN and RNN
We explore several advanced LSTM architectures. For Conv1D + BiLSTM, it combines 1-dimensional convolutional layers with bidirectional LSTM layers. The convolutional layer is used to extract features from the data [kimConvolutionalNeuralNetworks2014], while the bidirectional LSTM layer capture contextual information from both the past and future time steps [schusterBidirectionalRecurrentNeural1997]. For Conv1D + BiLSTM + Attention, the addition of an attention mechanism enables the model to focus on specific parts of the input data, improving performance by giving higher weights to more relevant parts [vaswaniAttentionAllYou2017]. This can help improve the accuracy of classification and facilitate further research on the importance of different phases for classification [salinasDistinguishingPlanetaryTransit2023].
@@ 3.1 CNN and RNN
In addition to the deep learning approaches, it is also important to consider the performance of classic machine learning methods for comparison. One such method is LightGBM, a gradient-boosting decision tree (GBDT) framework that uses tree-based learning algorithms [keLightGBMHighlyEfficienta].
3.2 Transformer
Transformer are highly effective in managing data dependencies, a critical aspect of time series analysis. This results in favorable outcomes [kitaevReformerEfficientTransformer2019, liuPYRAFORMERLOWCOMPLEXITYPYRAMIDAL2022]. Their multiple attention mechanisms are skilled in identifying the most relevant parts of input data [vaswaniAttentionAllYou2017]. By combining a 1-dimensional convolutional layer with a transformer encoder layer, we can effectively capture both global dependencies and local interactions in the data. This approach addresses the transformer architecture’s limitation in detectinglocal nuances.
@@ 3.2 Transformer
| Input Format | Model | Accuracy(All/T2CEP) | Macro F1-Score |
| Time-Series | Conv1D+Transformer | 85%/17% | 0.7 |
| LightGBM | 87%/25% | 0.71 | |
| BiLSTM+Attention | 93% | 0.76 | |
| Conv1D+BiGRU | 93% | 0.76 | |
| Conv1D+BiLSTM | 94% | 0.77 | |
| CWT Image | Efficientnet | 99% | 0.82 |
| Swin Transformer | 99%/83% | 0.96 | |
| Textual Time-Series | LLM-based model | 89% | 0.80 |
| Lightcurve Image | MLLM-based model | 95% | 0.94 |
| Transformed Audio | LALM-based model | 93% | 0.71 |

### v2

# Deep Learning and Methods Based on Large Language Models Stellar Light Curve Classification

@@ Abstract
Light curves serve as a valuable source of information on stellar formation and evolution. With the rapid advancement of machine learning techniques, they can be effectively processed to extract astronomical patterns and information. In this study, we present a comprehensive evaluation of models based on deep learning and large language models (LLMs) for the automatic classification of variable star light curves, using large datasets from the Kepler and K2 missions. Special emphasis is placed on Cepheids, RR Lyrae, and eclipsing binaries, examining the influence of observational cadence and phase distribution on classification precision. Employing automated deep learning optimization, we achieve striking performance 2 architectures: one that combines 1-dimensional convolution (Conv1D) with bidirectional long short-term memory (BiLSTM) and another called the Swin Transformer. These achieved accuracies of 94% and 99% respectively, with the latter demonstrating a notable 83% accuracy in discerning the elusive Type II Cepheids that comprise merely 0.02% of the total dataset. We unveil StarWhisper LightCurve (LC), a series of 3 LLM models based on an LLM, a multimodal large language model (MLLM), and a large audio language model (LALM). Each model is fine-tuned with strategic prompt engineering and customized training methods to explore the emergent abilities of these models for astronomical data. Remarkably, StarWhisper LC series models exhibit high accuracies around 90%, considerably reducing the need for explicit feature engineering, thereby paving the way for streamlined parallel data processing and the progression of multifaceted multimodal models in astronomical applications. The study furnishes 2 detailed catalogs illustrating the impacts of phase and sampling intervals on deep learning classification accuracy, showing that a substantial decrease of up to 14% in observation duration and 21% in sampling points can be realized without compromising accuracy by more than 10%.
Keywords: Computational astronomy, Variable stars, Large language model, Deep learning
Introduction
@@
The phenomenon of light variation is a crucial aspect of astrophysics and has long been studied in the field of time-domain astronomy. Cepheid variables, a special kind of variable star, serve a critical role as a universal standard candle, enabling us to measure the distance to clusters and galaxies with their period-luminosity relation. Other types of variable stars, including RR Lyrae (RR), Sct (DSCT), and Dor (GDOR) stars, are characterized by their unique absolute magnitudes, pulsation patterns, each exhibiting diverse period-luminosity relations. The classification of these pulsating stars has considerably enriched our understanding of their formation and evolution. Moreover, it has shed light on the structure and dynamics of binary, n-body systems, and our galaxy. [prsaKEPLERECLIPSINGBINARY2011, slawsonKEPLERECLIPSINGBINARY2011, matijevicKEPLERECLIPSINGBINARY2012, conroyKEPLERECLIPSINGBINARY2014a, conroyKeplerEclipsingBinary2014b, lacourseKeplerEclipsingBinary2015, kirkKEPLERECLIPSINGBINARY2016, abdul-masihKEPLERECLIPSINGBINARY2016a]
@@
Long periods of observations are crucial in order to understand the nature of variable stars. To carry out these observations, it is important to make wise use of telescope time, which is a valuable resource [garcia-piquerEfficientSchedulingAstronomical2017]. One strategy for doing so involves the use of simulations to create plans that optimally allocate available telescope time, considering overheads and other factors. Another strategy employs autonomous agents to optimize the observation of time-varying phenomena [saundersOptimalObservingAstronomical]. In this study, we aim to understand the importance tied to different phases of variables and the classification accuracy provided by varying cadences. This can help us to establish robust guidelines for future star gazing observations [wangTransferLearningApplied2023].
@@
A substantial amount of labeled data is required, in order to understand the significance of phase and cadence for different variable stars. In recent years, astronomy has entered the era of big data. For example, the Zwicky Transient Facility (ZTF) [bellmZwickyTransientFacility2014] at Palomar Observatory scans the sky every 2 days with a 1.2m telescope , generating approximately 1 TB of raw data each night [mahabalMachineLearningZwicky2019a]. The upcoming SiTian survey [liuSiTianProject2021], aims to monitor over 1,000 of sky every 30 minutes using fifty 1m class Schmidt telescopes, and is expected to produce around 140TB of processed data each night.
@@
As the volume of data grows, there is an increasing requirement for efficient and automated interpretation and analysis methods. Deep learning, a component of machine learning, has become a powerful tool for image and signal processing [lecunDeepLearning2015]. It can learn and extract features from data [krizhevskyImageNetClassificationDeep2012a, lecunDeepLearning2015], making it an ideal solution for managing large, complex datasets. Specifically, recurrent neural networks (RNNs) have shown great effectiveness in processing time series data [liptonCriticalReviewRecurrent2015], while convolutional neural networks (CNNs) have demonstrated their superiority in image processing tasks [krizhevskyImageNetClassificationDeep2012a, heDeepResidualLearning2015]. Moreover, the transformer architecture, with its attention mechanism [vaswaniAttentionAllYou2017], has shown remarkable performance across various applications [devlinBERTPretrainingDeep2019a, radfordImprovingLanguageUnderstanding, radfordLanguageModelsAre, brownLanguageModelsAre2020]. RNNs originally developed by [elmanFindingStructureTime1990], led to Simple RNNs, which were later enhanced by long short-term memory (LSTM, [hochreiterLongShortTermMemory1997]) to address long-term dependencies. These were further simplified by gated recurrent units (GRUs) [choPropertiesNeuralMachine2014] for computational efficiency. CNNs initially introduced for hand written digit recognition [lecunBackpropagationAppliedHandwritten1989], and have evolved with architectures such as AlexNet, VGGNet, and ResNet. Recent advancements include EfficientNet [tanEfficientNetRethinkingModel2020, tanEfficientNetV2SmallerModels2021], which optimizes depth, width, and resolution for model accuracy and efficiency.
@@
In recent years, transformer architecture has shown remarkable performance across a diverse range of applications.Initially introduced to address sequential data processing tasks [vaswaniAttentionAllYou2017]. Transformers use self-attention mechanisms to model relationships with sequences, enabling them to capture long-range dependencies and relationships between sequence elements. The success of the transformer in natural language processing has sparked interest in their potential application in other domains, such as computer vision and speech recognition. An example of a vision transformer is the Swin Transformer [liuSwinTransformerHierarchical2021, liuSwinTransformerV22022], which has optimized the computation of the attention mechanism and shown promising results in various benchmark datasets.
@@
Beyond the traditional reliance on extensive data for transfer learning, it is imperative to investigate the potential of large language models (LLMs) that incorporate both substantial datasets and large numbers of parameter. StarWhisper (available at https://github.com/Yu-Yang-Li/StarWhisper) is an LLM for astronomy that has strong astronomical ability and instruction-following ability and can complete a series of functions such as knowledge question answering, calling multimodal tools, and docking telescope control systems. The StarWhisper LC Series is initiated with the purpose of leveraging the experience gained from the prior training of the StarWhisper language model. It seeks to explore and discuss the potential harnessed from vast data to engender emergent properties in the analysis of light curve data. LLMs such as the Gemini 7B model have shown promise in adapting to new data types through fine-tuning with specific prompt templates [zhou2023fits13]. Multimodal large language models (MLLMs) such as the deepseek-vl-7b-chat, are adept at handling tasks involving image classification due to their extensive training on datasets containing chart data [tsai2019multimodal_6]. Large audio language models (LALMs), such as Qwen-audio,trained on audio datasets, exhibit exceptional performance in audio classification [yang2022voice2series_4].
@@
In this study, we perform a comprehensive evaluation of deep-learning and LLM-based models for the classification of variable star light curves, using the data from Kepler and K2 observations. The types of variable stars and the data pre-processing are presented in the section on data. In the section on models, we utilize various classification models, including LSTM, GRU, transformer, LightGBM, EfficientNet, the Swin Transformer and the StarWhisper LC series. Then we introduce training methods. We also present model performance and catalogs of the phase importance and sampling intervals in the section on results. A discussion is provided in the discussion section.
Materials and Methods
Data
Kepler and K2
@@
The Kepler spacecraft was launched in 2009 aiming to discover Earth like planets [doi:10.1126/science.1185402]. It was equipped with an optical telescope with a 95 cm aperture and a field of view. The advanced technology allowed Kepler to precisely track light curves from 200,000 different targets. This high precision resulted in the discovery of over 2,000 planets. For stars with V band magnitude between 13 mag and 14 mag, the precision was 100 parts per million (ppm), while for stars with V band magnitude between 9 mag to 10 mag, the precision was 10 ppm. Unfortunately, after 4 years, half of Kepler’s 4 reaction wheels failed, leading to the end of its primary mission. Yet, this marked the beginning of the K2 mission. The K2 mission used the transit method to detect changes in light along the ecliptic plane and created catalogs with photometric precision closely matching that of the original Kepler mission. The observations of K2 were controlled using the remaining reaction wheels and thrusters, with each campaign limited to 80 days.
@@
There are 2 types of light curves available: presearch data conditioning (PDC) and simple aperture photometry (SAP) light curves [Cleve2016KeplerIH]. SAP light curves preserve long-term trends, whereas PDC light curves, produced by the Kepler Operations Center, are corrected for systematic errors. Consequently, we opt for PDC light curves in our analysis. Prior to testing, we evaluated 2 data types with distinct temporal resolutions: long-cadence data, sampled every 30 minutes, and short-cadence data, sampled every minute. Due to the limited availability of short-cadence data, we decided not to use it and focused exclusively on long-cadence light curves. For each light curve, we eliminate quarter-to-quarter variations and normalize the flux to relative flux, following the methodologies outlined in [yangFlareCatalogFlare2019, hanStellarActivityCycles2021].
Variable stars
@@
Our training samples are similar to [wangTransferLearningApplied2023], including eclipsing binaries (EBs), RRs, DSCT, GDORs, and DSCT / GDOR hybrids (HYBs). The type II Cepheids (T2CEP) are included, in order to make a more universal sample. Table 1 lists our training samples and corresponding references. The training samples are seriously biased among different variable types. This bias often leads to decreased performance, especially in few-shot situations and small sample scenarios [kokolMachineLearningSmall2022, clemenconStatisticalLearningBiased2019]. However, [taniguchiMachineLearningModel2018] suggest that certain models may help overcome these challenges, although such cases seem to be the exception rather than the rule. These biased training samples offer a unique opportunity to study the algorithm’s dependence on such biases and to understand how this dependence affects the overall performance and accuracy when applied to astronomical data.
Pre-processing
@@
A pre-processing method, as described by [wangTransferLearningApplied2023], was adopted to manage and clean the light curves, with the aim of enhancing their features and expanding the training data. The lightcurves were segmented into 10-day intervals. Any segments with gaps exceeding 1 day were removed, while those with gaps less than 1 day were interpolated with a time sequence of 30-minute intervals.
| Label | Input Catalog | Final Samples | References |
| DSCT | 1,389 | 111,528 | [kholopovCombinedGeneralCatalogue1998], [durlevichListErrorsGCVS1994], [artyukhinaVizieROnlineData1996], [bradleyRESULTSSEARCHDOR2015], [balonaGaiaLuminositiesPulsating2018] |
| EB | 2,908 | 226,937 | [prsaKEPLERECLIPSINGBINARY2011], [slawsonKEPLERECLIPSINGBINARY2011], [matijevicKEPLERECLIPSINGBINARY2012], [conroyKEPLERECLIPSINGBINARY2014a] , [conroyKeplerEclipsingBinary2014b], [lacourseKeplerEclipsingBinary2015], [kirkKEPLERECLIPSINGBINARY2016], [abdul-masihKEPLERECLIPSINGBINARY2016] |
| GDOR | 941 | 65,786 | [bradleyRESULTSSEARCHDOR2015], [balonaGaiaLuminositiesPulsating2018] |
| HYB | 1,552 | 33,751 | [kholopovCombinedGeneralCatalogue1998], [durlevichListErrorsGCVS1994], [artyukhinaVizieROnlineData1996], [samus84thNameListVariable2021], [bradleyRESULTSSEARCHDOR2015], [balonaGaiaLuminositiesPulsating2018] |
| RR | 482 | 9,306 | [molnarGaiaDataRelease2018] |
| T2CEP | 3 | 94 | [kholopovCombinedGeneralCatalogue1998], [durlevichListErrorsGCVS1994], [artyukhinaVizieROnlineData1996], [samus84thNameListVariable2021], [molnarGaiaDataRelease2018] |
| Total | 7,275 | 447,402 |
@@
The input catalog column indicates the sources of data that we obtained from several catalogs. The final sample column represents the final sample size after applying all the preprocessing procedures.
@@
The table shows the period of the stars (in days) along with the variation of accuracy and time saved (percentages), for both phase importance and sampling research.
An alternative approach to time-series data classification involves the categorizing of images that are created from graphically represented light curves. This is achieved using transfer learning techniques. As highlighted in their result, the continuous wavelet transform (CWT) method has demonstrated superior results in imaging light curves.
The Morlet wavelet was chosen as the core function for the CWT, owing to its proven effectiveness in analyzing signals that display shifts in amplitude. This analysis resulted in a collection of images that provide valuable insights into the time-frequency characteristics of the signals. Importantly, these images facilitate the identification of patterns, trends, and anomalies, thus offering a comprehensive understanding of the signals.
Fig. 1 shows some variable stars from the training sample. The CWT image displays a time-frequency diagram, with the highest frequencies at the top of the image. T2CEP stars have the longest periods, so the top and middle parts of the CWT image are empty [percy2011Understanding]. RR stars pulsate with a very stable period of about 0.5 days and consistent amplitude, resulting in clear features at the top of the CWT images. HYB stars exhibit multiple types of variability, leading to features throughout the CWT images. DSCTs have multiple pulsation modes with periods ranging from 0.03 to 0.3 days [dup05DSCT], which are shorter than the pulsation periods of the GDORs, which range from 0.3 to 3 days. The CWT images of DSCTs show higher-frequency features compared to those of GDOR stars, and also shows different pulsation modes. For EBs, the light variations are caused by the eclipses of the 2 stars, resulting in a single clear feature in the CWT image. The periods of EBs vary, and some may have longer eclipse durations, represented by deeper straight bars.
Lomb-Scargle periodogram
We employ the Lomb-Scargle algorithm to extract the most significant periods complete light curves, which are used to study the phase importance of periodic variables. The Lomb-Scargle algorithm, a variant of the discrete fourier transform (DFT) developed by [1976Ap&SS..39..447L] and [scargleStudiesAstronomicalTime1982], has been specifically designed for unevenly sampled time-series data. It transforms a time series into a linear combination of sinusoidal waveforms, simplifying the conversion from the time domain to the frequency domain. We calculate the period by applying the Lightkurve Collaboration’s method [vanderplasUnderstandingLombScarglePeriodogram2018]. Some examples are shown in Figure 3.
@@
To assess the statistical significance of detected periods, we calculate the False Alarm Probability (FAP) for each period. The FAP is computed using the method provided by the LombScargle object in the Lightkurve library [2018ascl.soft12013L]. This method evaluates the probability of observing a signal of equal or greater power in random noise data. We adopt a threshold of FAP <0.001 to identify statistically significant periods. This criterion indicates that there is less than a 0.1% chance that a detected period could arise from random fluctuations in the data, thus providing a rigorous basis for distinguishing genuine periodic signals from noise.
Model construction
In order to classify variable stars from their light curves, we have explored various deep learning models. The basic architectures of these models are shown in Figure 4 and 5. Our model selection was guided by the specific characteristics of variable star light curves and the unique advantages offered by each architecture.
CNN and RNN models, like 1-dimensional convolutional layers (Conv1D) combined with bidirectional long short-term memory (BiLSTM), were selected due to their proficiency in capturing temporal dependencies and local characteristics in time series data [schusterBidirectionalRecurrentNeural1997, kimConvolutionalNeuralNetworks2014]. The bidirectional aspect of BiLSTM aids in identifying long-term dependencies, whereas Conv1D layers excel at extracting local features, thus enhancing the overall context. Transformer models, characterized by their self-attention mechanism [vaswaniAttentionAllYou2017], stand out in managing long sequences and capturing global dependencies, making them ideally suited for analyzing variable star data with intricate periodic patterns. In our image-based methods employing CWT, we utilized the Swin Transformer [liuSwinTransformerHierarchical2021, liuSwinTransformerV22022] and EfficientNet [tanEfficientNetRethinkingModel2020, tanEfficientNetV2SmallerModels2021]. The hierarchical windowed attention mechanism of the Swin Transformer is adept at capturing multi-scale features, which is particularly beneficial in few-shot learning contexts. Meanwhile, EfficientNet achieves an optimal balance between network depth, width, and resolution to ensure efficient feature extraction. Furthermore, we investigated the application of large language models via the StarWhisper LC series (LLM, MLLM, LALM), harnessing their robust generalization abilities and adaptability to novel tasks in time series data classification.
CNNs and RNNs
We explore several advanced LSTM architectures. Conv1D + BiLSTM combines 1-dimensional convolutional layers with bidirectional LSTM layers. The convolutional layer is used to extract features from the data [kimConvolutionalNeuralNetworks2014], while the bidirectional LSTM layer captures contextual information from both past and future time steps [schusterBidirectionalRecurrentNeural1997]. The attention mechanism of the Conv1D + BiLSTM + Attention model enables it to focus on specific parts of the input data, improving performance by giving higher weights to more relevant parts [vaswaniAttentionAllYou2017]. This can improve the accuracy of classification and facilitate further research on the importance of different phases for classification [salinasDistinguishingPlanetaryTransit2023].
@@
In addition to deep learning approaches, it is also important to consider the performance of classic machine learning methods for comparison. One such method is LightGBM, a gradient-boosting decision tree (GBDT) framework that uses tree-based learning algorithms [keLightGBMHighlyEfficienta].
Transformers
Transformers are highly effective in managing data dependencies, a critical aspect of time series analysis. This results in favorable outcomes [kitaevReformerEfficientTransformer2019, liuPYRAFORMERLOWCOMPLEXITYPYRAMIDAL2022]. Their multiple attention mechanisms are skilled in identifying the most relevant parts of input data [vaswaniAttentionAllYou2017]. By combining a 1-dimensional convolutional layer with a transformer encoder layer, we can effectively capture both global dependencies and local interactions in the data. This approach addresses the transformer architecture’s limitation in detecting local nuances.
@@
| Input Format | Model | Accuracy (All/T2CEP) | Precision | Recall | Macro F1- Score |
| Time-Series | Conv1D + Transformer | 85%/17% | 0.69 | 0.76 | 0.70 |
| LightGBM | 87%/25% | 0.70 | 0.81 | 0.71 | |
| BiLSTM + Attention | 93% | 0.91 | 0.92 | 0.76 | |
| Conv1D + BiGRU | 93% | 0.91 | 0.91 | 0.76 | |
| Conv1D + BiLSTM | 94% | 0.92 | 0.94 | 0.77 | |
| CWT Image | EfficientNet | 99% | 0.98 | 0.99 | 0.82 |
| Swin Transformer | 99%/83% | 0.96 | 0.95 | 0.96 | |
| Textual Time-Series | LLM-based model | 89% | 0.77 | 0.87 | 0.80 |
| Light Curve Image | MLLM-based model | 95% | 0.93 | 0.95 | 0.94 |
| Transformed Audio | LALM-based model | 93% | 0.88 | 0.90 | 0.71 |